3. Virtual Memory (20 points) An ISA supports an 8
Answer #1

3. A TLB misses /page faults:

LRU

TLB(VPN)

PHYSICAL PAGE #

          ?

3

5

?

4

3

?

7

1

3

3

5

7

7

1

1

1

3

4

4

5

10

10

1

10

10

1

7

7

3

11

11

5

11

11

5

7

7

3

8

8

1

10

10

5

8

8

1

7

7

3

12

12

5

9

9

1

                                                                                                   

8 , 10, 8 , 1, 12 , 9

TLB Hit Ratio: total repeated sequence /Total Number of Sequence= 4/16=0.22 (22.22% in terms of percentage)

TLB Hits: 3, 7, 4

Page Faults: 1,4,10, 11, 8, 12, 9

3. B Final TLB:

LRU

TLB(VPN)

PHYSICAL PAGE #

7

7

3

12

12

5

9

9

1

3. C Physical Frame Contents are

FRAME(PHYSICAL PAGE ) NUMBER

FRAME CONTENTS

0

Empty

1

9

2

1

3

7

4

Empty

5

12

6

Empty

7

Page Table